An object is observed for 5s. The 1st second the o


Answer the pictured question in its entirety and graph. Thanks!!
Answer #1

apply the formula for accleration a from S = 0.5at^2

where s is distance travelled

so

during 1st second, accleration a1 = 2* 6/1*1 = 12 m/s^2

during 2nd second, accleration a2 = 2 * 6/1*1 = 12m/s^2

for third second a3 = 2 * 6/1 = 12 m/s^2

even for 4 and 5th seconds a4 = a5 = 12 m/s^2